OpenStudy (wyattp17):

6 square root of nine and ur trying to simplifying radicals

OpenStudy (wyattp17):



OpenStudy (luigi0210):

Here: \[\LARGE 6\sqrt{9}=6\sqrt{3*3}\]

OpenStudy (luigi0210):

and \[\LARGE \sqrt{x*x}=x\]

OpenStudy (wyattp17):

i dont get it still @Luigi0210

OpenStudy (luigi0210):

There's two 3's in there so it comes out.. and you are left with: \[\LARGE 6*3\]

OpenStudy (wyattp17):

what @Luigi0210

OpenStudy (wyattp17):

ohh @Luigi0210

OpenStudy (anonymous):